OpenStudy (anonymous):

Josh types the five entries in the bibliography of his term paper in random order , forgetting that thet should be in alphabetical oder by author ,What is the probability that he actually typed them in a alphabetical order?

OpenStudy (anonymous):

\[\frac{1}{5!}\]

OpenStudy (anonymous):

probability first one is right is \[\frac{1}{5}\] then probability second one is right given the first one is right is \[\frac{1}{4}\] etc. then multiply since they all have to be right

OpenStudy (anonymous):

a more prosaic way to think about it is that there are 5! different ways to arrange, and only one is correct. same thing really